OpenStudy (anonymous):

If x is a number that is 7 more than y, how do you express x as a function of y?

OpenStudy (condor34):

f(x)=x+7?

OpenStudy (anonymous):

x = f(y) = 7 + x x = f(y) = 7 + y y = f(x) = 7x y = f(x) = y – 7 y = f(x) = 7 + y options

OpenStudy (jdoe0001):

so, whatever "y" is, "x" is 7 more than that thus x = y + 7

OpenStudy (jdoe0001):

y+7 or 7+y the commutative property works the same anyhow

OpenStudy (anonymous):

so it's either 2 or 5

OpenStudy (mertsj):

x=7+y so x=f(y)=7+y